# Retrieval Augmented Diffusion Model for Structure-informed Antibody Design and Optimization

### v1

@@ 5.1 Antibody CDR Sequence inverse folding
Baselines For traditional methods, we simulated a method of grafting using CDR-like data in the process of rational antibody design. Specifically, we directly graft the retrieved top-1 CDR-like fragment onto the antibody framework, termed Grafting. For deep learning methods, we selected a series of state-of-the-art protein inverse folding models for comparison with our work, including ProteinMPNN (Dauparas et al., 2022), a model that utilizes message passing neural network to design sequences with a fixed protein backbone; ESM-IF (Hsu et al., 2022), a protein inverse folding model that trained on millions of predicted structures; Diffab-fix (Luo et al., 2022), which can fix the backbone structure and iteratively generate candidate sequences from pure noise in sequence space using diffusion; AbMPNN (Dreyer et al., 2023), a model fine-tuned ProteinMPNN on antibody sequence and structure. Because it is not open-sourced, we evaluate it on its own test set. For more baseline details, please refer to Appendix A.5.
@@
Results As shown in Table 1, RADAb outperforms state-of-the-art methods in each metric and at each CDR region. In particular, in the highly variable and specific CDR-H3 region (Shirai et al., 1999; Raybould et al., 2019), our method achieved a great improvement in AAR compared to best-performing methods Diffab-fix and AbMPNN. The evaluation results indicate that the retrieval-augmented method, by introducing structurally similar homologous sequences, has improved the accuracy, consistency, and rationality of the model’s generation.
In addition, CDR-H3 exhibits significant variability in length, sequence, and structure. Typically, deep learning models show decreased performance when generating longer CDR-H3 sequences (Luo et al., 2022; Høie et al., 2024). Therefore, we selected a subset of the test set with CDR-H3 lengths longer than 14 to evaluate the generation performance. As shown in Table 3, while the generation performance of all methods declines to some extent, our method demonstrates consistency and significantly outperforms the others, with a larger margin of improvement.
@@ 5.2 Antibody functionality optimization
In this section, we focus on the evolution of antibody sequences and evaluate whether the structure of the evolved sequence has greater functionality compared to the structure of the folded original sequence. To this end, we first fold the designed CDR-H3 sequences with framework sequences and the original real antibody sequences into complete protein structures using ABodyBuilder2. Then, we use FastRelax and InterfaceAnalyzer in PyRosetta (Alford et al., 2017) to relax the structure and calculate the binding energy $\Delta$G of the antibody-antigen complex.
Metrics We use various metrics to evaluate the efficacy and functionality of our designed antibodies: (1) $\Delta$$\Delta$G: This metric represents the difference in binding energy between the complex with the designed CDR folded into the structure and the original complex binding energy. (2) $\Delta$$\Delta$G-seq: This metric measures the difference in binding energy between the complex with the designed CDR sequence folded into the structure and the binding energy of the original antibody sequence folded into the structure. It aims to eliminate errors introduced by the folding tool, allowing for a direct comparison of sequence functionality. (3) IMP-seq: This metric indicates the percentage of designed CDR sequences folded into the structure with a lower (better) binding energy than the original antibody sequences folded into the structure.
@@
To further demonstrate the optimization of antibody sequence functionality by RADAb, we select a specific antigen-antibody complex from the test set (A neutralizing MAb targeting the receptor-binding domain of SARS-CoV-2, PDB: 7d6i). We generate 50 sequences for CDR-H3 and calculate the binding energy $\Delta$G of the folded structures. Among these, 68% of the samples exhibited lower $\Delta$G compared to the original complex. As shown in Figure 3, we select two samples as examples. Although they do not achieve the highest AAR, they demonstrate better binding affinity compared to the native structure.
@@ A.3 Implementation of structural retrieval
Assume that the coordinates of residue $j$ are aligned with the central residue of $\mathcal{X}$, and then compute the RMSD of $\mathcal{X}$ when aligned onto $\tau$. If the input contains discontinuous multiple structures, cRMSD will be the cumulative RMSD of these structures. MaxA, maxB, and maxC are three different upper-bound thresholds. These thresholds are selected to improve the speed and accuracy of the retrieval algorithm (For detailed proof, please refer to MASTER (Zhou & Grigoryan, 2015)).
@@ A.4 Implementation of CDR-like database constructing
To eliminate the computational overhead caused by structural retrieval during the model’s training and inferencing, we followed previous work (Aguilar Rangel et al., 2022) and initially executed the retrieval algorithm on all CDR structures of all antibodies to construct a CDR-like database.
@@ A.5.1 Traditional methods
Rosetta-Fixbb (Adolf-Bryfogle et al., 2018) Rosetta-Fixbb can use energy functions for antibody CDR sequence design. Since DiffAb has already been proven to outperform it (Luo et al., 2022; Wu & Li, 2023) on sequence design task, we did not conduct additional comparisons.
@@ A.5.2 Deep learning methods
ProteinMPNN (Dauparas et al., 2022) ProteinMPNN is a deep learning framework for protein sequence inverse folding. It leverages a message passing neural network to model the complex relationships between amino acids in a protein structure. We use the antibody’s backbone structure as input and keep the sequences outside the CDR regions to be designed fixed. We design sequences for each CDR region separately. The sampling temperature is set to the default value of 0.1.
Esm-IF1 (Hsu et al., 2022) Esm-IF1 is a protein sequence inverse folding model trained on millions of AlphaFold2 predicted structures. We use the antibody’s backbone structure as input and keep the sequences outside the CDR regions to be designed fixed. We design sequences for each CDR region separately. The sampling temperature is set to the value of 0.2.
Diffab-fix (Luo et al., 2022) Diffab is a diffusion model that can design sequences of CDR region with a fixed CDR backbone. It takes antigen-antibody framework context as condition to design CDR sequence. For a fair comparison, we retrained it with the default training configuration fixbb.yml.
AbMPNN (Dreyer et al., 2023) AbMPNN is fine-tuned by antibody structure data and predicted OAS (Observed Antibody Space ) structure data. Its model architecture is consistent with ProteinMPNN but achieves better performance in antibody inverse folding. We use the antibody’s backbone structure as input and keep the sequences outside the CDR regions to be designed fixed. However, it is not open-sourced yet, so we evaluate it on its own test set. We design sequences for each CDR region separately. The sampling temperature is set to the default value of 0.1.
@@ A.6 Experiment on CDR-H3’s length
Appendix B Restrict the evolution space by structural retrieval
The search space for protein sequences is vast. If the length of the sequence to be designed is L, the search space is $20^{L}$. General deep learning methods have already significantly reduced the sequence search space, generating sequences of much higher quality than random sequences. However, it is still challenging to find mutation spaces with high fitness. To restrict this space, we first fix the backbone structure of the CDR region, limiting the space to a specific area. Subsequently, we add more constraints to the mutation space by retrieving the PDB database for motifs similar to the CDR backbone structure and taking them as the model’s conditions, as shown in Figure 2 C.
Appendix C Structure reconstruction
To reconstruct the antibody structure, we use ABodyBuilder2 (Abanades et al., 2023), a deep learning model capable of predicting antibody light chain-heavy chain complexes. It is significantly faster than AlphaFold2 and offers higher prediction accuracy. We insert the designed CDR sequences into the antibody framework sequence, input it into ABodyBuilder2 to fold, and use OpenMM relax to obtain the structure corresponding to the new CDR sequences. Subsequently, we align the structure to the real antibody framework. Finally, we use the fastrelax function in PyRosetta (Alford et al., 2017), with the score function set to ref2015 and max iteration set to 1000, to relax the structure.
Appendix D Training and inference algorithm
@@ D.2 Sampling algorithm
Appendix E Case Study
We select a portion of the optimized antibodies in Figure S2. They achieved lower binding energy compared to the original antibody structures. Various examples demonstrate that redesigning antibody sequences in the CDR regions by fixing the backbone structure can produce antibodies with lower antigen-binding energy, even if the designed sequences do not have high similarity to natural antibodies. This method is a successful approach to optimizing antibody functionality.
@@ Appendix F Limitations
Quality of retrieved data Due to the inherent limitations of the MASTER algorithms, the retrieved linear motifs may have structural issues. Despite careful screening and filtering, a tiny portion of the data might have lengths that differ from the original CDR loops or may even become discontinuous due to missing residues. These exceptional cases may have a negative impact on our model. We hope that advances in structural retrieval and improvements in alignment will jointly address this issue.
Computational Cost When constructing the CDR-like fragments database, searching the entire PDB database using all CDR structures from the SabDab database takes approximately 100 hours. Additionally, the ESM2 encoding used to capture antibody sequence evolution information and the axial attention focused on local CDR in the denoising network require more computational resources than typical diffusion methods.

### v2

@@ 5.1 Antibody CDR Sequence inverse folding
Baselines For traditional methods, we simulated a method of grafting using CDR-like data in the process of rational antibody design. Specifically, we directly graft the retrieved top-1 CDR-like fragment onto the antibody framework, termed Grafting. For deep learning methods, we selected a series of state-of-the-art protein inverse folding models for comparison with our work, including ProteinMPNN (dauparas2022robust), a model that utilizes message passing neural network to design sequences with a fixed protein backbone; ESM-IF (hsu2022learning), a protein inverse folding model that trained on millions of predicted structures; Diffab-fix (luo2022antigen), which can fix the backbone structure and iteratively generate candidate sequences from pure noise in sequence space using diffusion; AbMPNN (dreyer2023inverse), a model fine-tuned ProteinMPNN on antibody sequence and structure. Because it is not open-sourced, we evaluate it on its own test set. For more baseline details, please refer to Appendix A.5.
@@
Results As shown in Table 1, RADAb outperforms state-of-the-art methods in each metric and at each CDR region. In particular, in the highly variable and specific CDR-H3 region (shirai1999h3; raybould2019five), our method achieved a great improvement in AAR compared to best-performing methods Diffab-fix and AbMPNN. The evaluation results indicate that the retrieval-augmented method, by introducing structurally similar homologous sequences, has improved the accuracy, consistency, and rationality of the model’s generation.
In addition, CDR-H3 exhibits significant variability in length, sequence, and structure. Typically, deep learning models show decreased performance when generating longer CDR-H3 sequences (luo2022antigen; hoie2024antifold). Therefore, we selected a subset of the test set with CDR-H3 lengths longer than 14 to evaluate the generation performance. As shown in Table 3, while the generation performance of all methods declines to some extent, our method demonstrates consistency and significantly outperforms the others, with a larger margin of improvement.
@@ 5.2 Antibody functionality optimization
[…]
@@
[…]
@@ 6 Conclusion and Future works
Reproducibility statement
The code is avalibale at https://github.com/GENTEL-lab/RADAb
Acknowledgments
This study has been supported by the National Natural Science Foundation of China [62041209], Natural Science Foundation of Shanghai [24ZR1440600], the Science and Technology Commission of Shanghai Municipality [24510714300].
@@ A.3 Implementation of structural retrieval
Assume that the coordinates of residue $j$ are aligned with the central residue of $\mathcal{X}$, and then compute the RMSD of $\mathcal{X}$ when aligned onto $\tau$. If the input contains discontinuous multiple structures, cRMSD will be the cumulative RMSD of these structures. MaxA, maxB, and maxC are three different upper-bound thresholds. These thresholds are selected to improve the speed and accuracy of the retrieval algorithm (For detailed proof, please refer to MASTER (zhou2015rapid)).
@@ A.4 Implementation of CDR-like database constructing
To eliminate the computational overhead caused by structural retrieval during the model’s training and inferencing, we followed previous work (aguilar2022fragment) and initially executed the retrieval algorithm on all CDR structures of all antibodies to construct a CDR-like database.
@@ A.5.1 Traditional methods
Rosetta-Fixbb (adolf2018rosettaantibodydesign) Rosetta-Fixbb can use energy functions for antibody CDR sequence design. Since DiffAb has already been proven to outperform it (luo2022antigen; wuHTP) on sequence design task, we did not conduct additional comparisons.
@@ A.5.2 Deep learning methods
ProteinMPNN (dauparas2022robust) ProteinMPNN is a deep learning framework for protein sequence inverse folding. It leverages a message passing neural network to model the complex relationships between amino acids in a protein structure. We use the antibody’s backbone structure as input and keep the sequences outside the CDR regions to be designed fixed. We design sequences for each CDR region separately. The sampling temperature is set to the default value of 0.1.
Esm-IF1 (hsu2022learning) Esm-IF1 is a protein sequence inverse folding model trained on millions of AlphaFold2 predicted structures. We use the antibody’s backbone structure as input and keep the sequences outside the CDR regions to be designed fixed. We design sequences for each CDR region separately. The sampling temperature is set to the value of 0.2.
Diffab-fix (luo2022antigen) Diffab is a diffusion model that can design sequences of CDR region with a fixed CDR backbone. It takes antigen-antibody framework context as condition to design CDR sequence. For a fair comparison, we retrained it with the default training configuration fixbb.yml.
AbMPNN (dreyer2023inverse) AbMPNN is fine-tuned by antibody structure data and predicted OAS (Observed Antibody Space ) structure data. Its model architecture is consistent with ProteinMPNN but achieves better performance in antibody inverse folding. We use the antibody’s backbone structure as input and keep the sequences outside the CDR regions to be designed fixed. However, it is not open-sourced yet, so we evaluate it on its own test set. We design sequences for each CDR region separately. The sampling temperature is set to the default value of 0.1.
@@ A.6 Experiment on CDR-H3’s length
A.7 Detailed results of antibody functionality optimization
To further evaluate the model’s performance in optimizing antibody functionality, we additionally assessed the $\Delta\Delta G_{seq}$ of the top-1 to top-3 structures generated by the model. The results are shown in Table 5, which further demonstrate that the antibodies optimized by our model exhibit improved functionality.
[…]
Grafting 135.17 40.22 32.69 - - -
ProteinMPNN 127.14 24.72 35.51 -54.63 -45.58 -35.70
ESM-IF1 162.09 42.28 33.33 -62.65 -48.81 -33.86
Diffab-fix 116.36 14.05 34.52 -62.51 -54.80 -46.13
RADAb 109.16 7.06 37.30 -69.30 -55.95 -45.96
[…]
Appendix B Structure reconstruction
To reconstruct the antibody structure, we use ABodyBuilder2 (abanades2023immunebuilder), a deep learning model capable of predicting antibody light chain-heavy chain complexes. It is significantly faster than AlphaFold2 and offers higher prediction accuracy. We insert the designed CDR sequences into the antibody framework sequence, input it into ABodyBuilder2 to fold, and use OpenMM relax to obtain the structure corresponding to the new CDR sequences. Subsequently, we align the structure to the real antibody framework. Finally, we use the fastrelax function in PyRosetta (alford2017rosetta), with the score function set to ref2015 and max iteration set to 1000, to relax the structure.
Appendix C Training and inference algorithm
@@
Appendix D Case Study
We select a portion of the optimized antibodies in Figure S2. They achieved lower binding energy compared to the original antibody structures.
@@ Appendix D Case Study
Appendix E The motivation for only considering sequence design
Based on our observations, inverse folding represents a more practical scenario. Current structure-sequence co-design methods typically involve masking the CDR while retaining the presence of an antibody framework backbone, which represents a relatively uncommon use case. In most practical scenarios, we either have access to the full complex structure of the template antibody and the antigen, allowing us to perform inverse folding, or we lack a template molecule entirely, necessitating full atom and de novo design. This is also why researches from pharmaceutical companies and efforts involving in vitro experiments on antibody loop regions tend to focus more on designing antibodies through inverse folding, as evidenced by several recent studies (shanehsazzadeh2023vitro; shanehsazzadeh2023unlocking; freyprotein; hoie2024antifold; shanker2024unsupervised). This rationale underpins our decision to concentrate solely on this aspect in our work.
Another reason is that performing sequence-structure co-design while adhering to our retrieval-based approach would risk data leakage. This is because our retrieval process relies on the known structure of the CDR region, which is only possible when the CDR backbone is already known. At the same time, we recognize that epitope-specific full antibody de novo design, including full-atom design, is highly valuable. We are actively developing retrieval systems for PPI (Protein-Protein Interaction) retrieval to avoid potential data leakage and enhance our model.
@@ Appendix F Limitations
Due to the inherent limitations of the MASTER algorithms, the retrieved linear motifs may have structural issues. Despite careful screening and filtering, a tiny portion of the data might have lengths that differ from the original CDR loops or may even become discontinuous due to missing residues. These exceptional cases may have a negative impact on our model. We hope that advances in structural retrieval and improvements in alignment will jointly address this issue.
When constructing the CDR-like fragments database, searching the entire PDB database using all CDR structures from the SabDab database takes approximately 100 hours. Additionally, the ESM2 encoding used to capture antibody sequence evolution information and the axial attention focused on local CDR in the denoising network require more computational resources than typical diffusion methods.
It is important to note that applying current retrieval mechanism to structure and sequence co-design poses a significant risk of data leakage. Specifically, the retrieval process inherently relies on using the CDR structure information as a query, which practically transforms the task into one resembling inverse folding. However, one potential approach is to incorporate PPI (Protein-Protein Interaction) retrieval, and we are currently experimenting with MASIF(gainza2020deciphering; gainza2023novo) for this purpose. This will be explored further as part of our future work.